Let X be uniform on [0, 1], and let Y be exponential with rate λ, so that P(Y ≥ t) = e ^(−λt ) t ≥ 0 and 1 if t < 0

Assume that X and Y are independent, and define W = X + 2Y .

a) For any w ≥ 0 and x ∈ [0, 1], compute P(W ≥ w|X = x)

b) By undoing the conditioning on X, use the result from part (a) to compute the cdf of W.
Answer #1

x~ Unifolm (10, 13) yn Exponential (at) @ P(W zw/x=x) = P(x+2yZw/x= x) = P(x+2yzw) (this is valid because =P(2yZW-x) = P(yz W(6) We can write P(wz w) = P(wzw/x=x) fy() dx. where fy(x) is the probability density function of x <x > P(WZW) = P(wzw/x=x)P(W=w)= | P(w z w/ x=x) dx. - (W-x)/2 dx if w> = edula letele dx. - edula ex12 | - edula e diz) lo = endula ac etia_i). = 2